# footnotetext: * Equal Contribution

### v1

@@ Abstract
Monocular dynamic reconstruction is a challenging and long-standing vision problem due to the highly ill-posed nature of the task. Existing approaches are limited in that they either depend on templates, are effective only in quasi-static scenes, or fail to model 3D motion explicitly. In this work, we introduce a method capable of reconstructing generic dynamic scenes, featuring explicit, full-sequence-long 3D motion, from casually captured monocular videos. We tackle the under-constrained nature of the problem with two key insights: First, we exploit the low-dimensional structure of 3D motion by representing scene motion with a compact set of $\mathbb{SE}(3)$ motion bases. Each point’s motion is expressed as a linear combination of these bases, facilitating soft decomposition of the scene into multiple rigidly-moving groups. Second, we utilize a comprehensive set of data-driven priors, including monocular depth maps and long-range 2D tracks, and devise a method to effectively consolidate these noisy supervisory signals, resulting in a globally consistent representation of the dynamic scene. Experiments show that our method achieves state-of-the-art performance for both long-range 3D/2D motion estimation and novel view synthesis on dynamic scenes. Project Page: shape-of-motion.github.io
Keywords:
Long-range 3D motion tracking Monocular dynamic novel view synthesis 4D Reconstruction
Figure 1: Shape of Motion. Our approach enables joint 3D long-range tracking and novel view synthesis from a monocular video of a complex dynamic scene. Here, we demonstrate our ability to render moving scene elements at fixed viewpoint with different moments in time. Additionally, we visualize the estimated long-range 3D motion as colorful trajectories. These trajectories reveal distinct geometric patterns that encapsulate the movement of each point through 3D space and time, which leads us to the term “Shape of Motion”.
@@ 1 Introduction
Reconstructing the persistent geometry and their 3D motion across a video is crucial for understanding and interacting with the underlying physical world. While recent years have seen impressive progress in modeling static 3D scenes [61, 39], recovering the geometry and motion of complex dynamic 3D scenes, especially from a single video, remains an open challenge. A number of prior dynamic reconstruction and novel view synthesis approaches have attempted to tackle this problem. However, most methods rely on synchronized multi-view videos [8, 19, 19, 58, 99] or additional LIDAR/depth sensors [24, 91, 55, 68, 97]. Recent monocular approaches can operate on regular dynamic videos, but they typically model 3D scene motion as short-range scene flow between consecutive times [51, 52, 20] or deformation fields that maps between canonical and view space [64, 65, 108], failing to capture 3D motion trajectories persistent over a video.
The longstanding challenge for more general in-the-wild videos lies in the poorly constrained nature of the reconstruction problem. In this work, we tackle this challenge with two key insights. The first is that, while the image space dynamics can be complex and discontinuous, the underlying 3D motion is a composition of continuous simple rigid motions. Our second insight is that data-driven priors provide complementary, though noisy cues, that aggregate well into a globally coherent representation of the 3D scene geometry and motion.
Motivated by these two insights, we represent the dynamic scene as a set of persistent 3D Gaussians, and represent their motion across the video in terms of a compact set of shared $\mathbb{SE}(3)$ motion bases. Unlike traditional scene flow, which computes 3D correspondence between consecutive frames, our representation recovers a persistent 3D trajectory over the whole video, enabling long-range 3D tracking. As the 3D trajectories produced by our method capture the geometric patterns that trace each point’s movement through 3D space and time, we refer to our approach “Shape of Motion”, as shown in Figure 1. We show how to fit our explicit scene representation to a general video in-the-wild, by fusing together complementary cues from two main sources: monocular depth estimates per-frame, and 2D track estimates across frames. We conduct extensive evaluations on both synthetic and real-world dynamic video datasets, and show that our proposed approach significantly outperforms prior monocular dynamic novel view synthesis methods and 3D tracking baselines in both long-range 2D and 3D tracking accuracy. Moreover, we achieve state-of-the-art novel view synthesis quality among all existing methods. In summary, our key contributions include: (1) A new dynamic scene representation enabling both real-time novel view synthesis and globally consistent 3D tracking for any point at any time. (2) A carefully designed framework that optimizes the representation on posed monocular video by leveraging physical motion priors and data-driven priors.
@@ 2 Related Work
Monocular 3D long-range tracking has not been explored broadly in the literature, but there exist many approaches to perform tracking in 2D image space. A typical way for determining 2D point correspondences relies on optical flows. This involves estimating dense motion fields between image pairs [27, 4, 57, 9, 6, 7, 85, 15, 28, 36, 86, 31, 28, 101, 37, 35, 71, 79]. While effective for consecutive frames, accurate long-term tracking in videos remains a challenge with optical flow methods. Sparse keypoint matching methods can enable long trajectory generation [2, 73, 56, 12, 53], but these methods are primarily intended for sparse 3D reconstruction. Long-range 2D trajectory estimation for arbitrary points has been explored in earlier works, which relied on hand-crafted priors to generate motion trajectories [72, 3, 75, 76, 80, 92]. Recently, there has been a resurgence of interest in this problem, with several works showcasing impressive long-term 2D tracking results on challenging, in-the-wild videos. These approaches employ either test-time optimization where models consolidate noisy short-range motion estimates into a global representation for producing long-term correspondences [94, 62], or data-driven strategies [25, 38, 14], where neural networks learn long-term correspondence estimates from synthetic data [13, 112]. While these methods effectively track any 2D point throughout a video, they lack the knowledge of underlying 3D scene geometry and motions.
Scene flow or 3D motion trajectory is a common representation to model 3D scene motion and point correspondences. Most prior work estimates scene flow directly from Lidar point clouds [24, 91, 55, 68, 97] or RGBD images [84, 88, 34, 69, 89, 60]. In monocular setting, a few recent works proposed to estimate 3D motions through self-supervised learning or test-time optimization strategies [103, 30, 51, 52, 102, 104], but these approaches either focus on a single object, require template priors, or only produce short-range motion correspondences. In contrast, our method does not rely on template priors and is capable of producing long-range 3D trajectories, making it suitable for modeling complex scenes with multiple moving objects.
2.0.2 Dynamic Reconstruction and View Synthesis.
@@
Our work also relates to dynamic 3D scene reconstruction and novel view synthesis problems. In non-rigid reconstruction, early methods often required RGBD sensors [113, 63, 5, 16, 32] or strong hand-crafted priors [74, 45, 70]. Recent work has demonstrated progress toward the reconstruction of dynamic scenes in the wild by integrating monocular depth priors [50, 111, 59, 43, 110]. Recently, Neural Radiance Fields (NeRF)[61] and Gaussian Splat [40] based approaches have achieved state-of-the-art results. Most of these methods [1, 19, 8, 48, 19, 10, 81, 83, 93, 49] require simultaneous multi-view video observations or predefined templates [33, 47, 98] for high-quality novel view outputs. Template-free monocular approaches model dynamic scenes with different types of representations such as video depth maps [109], time-dependent NeRFs [17, 51, 67, 64, 90, 65, 100, 52], and dynamic 3D Gaussians [99, 18, 107, 108]. While significant progress has been made, as DyCheck [21] pointed out, many approaches focus on scenarios with camera teleportation [44, 108] or quasi-static scenes, which do not represent real-world monocular videos. In this work, we focus on modeling casual videos captured by a single camera, a more practical and challenging setup.
@@
Figure 2: System Overview. Given a single RGB video sequence with known camera poses, along with monocular depth maps and 2D tracks computed from off-the-shelf models [106, 14] as input, we optimize a dynamic scene representation as a set of persistent 3D Gaussians that translate and rotate over time. To capture the low-dimensional nature of scene motion, we model the motion with a set of compact $\mathbb{SE}(3)$ motion bases shared across all scene elements. Each 3D Gaussian’s motion is represented as a linear combination of these global $\mathbb{SE}(3)$ motion bases, weighted by motion coefficients specific to each Gaussian. We supervise our scene representation (canonical 3D Gaussian parameters, per-Gaussian motion coefficients, and global motion bases) by comparing the rendered outputs (RGB, depths and 2D tracks) with the corresponding input signals. This results in a dynamic 3D representation of the scene with explicit long-range 3D scene motion.
@@ 3 Method
Our method takes as input a sequence of $T$ video frames $\{I_{t}\in\mathbb{R}^{H\times W\times 3}\}$ of a dynamic scene, the camera intrinsics $\mathbf{K}_{t}\in\mathbb{R}^{3\times 3}$, and world-to-camera extrinsics $\mathbf{E}_{t}\in\mathbb{SE}(3)$ of each input frame $I_{t}$. From these inputs, we aim to recover the geometry of the entire dynamic scene and the full-length 3D motion trajectory of every point in the scene. Unlike most prior dynamic NeRFs methods [50, 52, 21, 100, 90] which render the scene contents through volumetric ray casting and represent the motion implicitly at fixed 3D locations, we model the dense scene elements as a set of canonical 3D Gaussians, but allow them to translate and rotate over entire video through full-length motion trajectories. We adopt explicit point-based representation because it simultaneously allows for both (1) high-fidelity rendering in real-time and (2) full-length 3D tracking of any surface point from any input time.
Optimizing an explicit representation of dynamic 3D Gaussians from a single video is severely ill-posed – at each point in time, the moving subjects in the scene are observed at those poses from only a single viewpoint. In order to overcome this ambiguity, we make two insights: First, we note that, while the projected 2D dynamics might be complex in the video, the underlying 3D motion in the scene is low-dimensional, and composed of simpler units of rigid motion. Second, powerful data-driven priors, namely monocular depth estimates and long-range 2D tracks, provide complementary but noisy signals of the underlying 3D scene. We propose a system that allows us to fuse these noisy estimates together into a globally conherent representation of both the scene geometry and motion.
The following sections introduce our framework. We represent the scene contents as a globally consistent set of 3D Gaussians that live in a canonical space (Section 3.1). To recover consistent motion trajectories, we bind the 3D Gaussians to a compact and low-dimensional motion parameterization. In particular, we represent the full-length motion trajectories of each dynamic scene element with a set of compact and low-dimensional $\mathbb{SE}(3)$ motion bases (Section 3.2). Finally, we fit these motion bases to the input video frames, constraining our optimization with structural priors and data-driven priors (Section 3.3). We show a schematic of our pipeline in Figure 2.
@@ 3.1 Preliminaries: 3D Gaussian Splatting
We represent the appearance and geometry of dynamic scene contents with a global set of 3D Gaussians, an explicit and expressive differentiable scene representation [39] for efficient optimization and rendering. We define parameters of each 3D Gaussian in the canonical frame $t_{0}$ as $\vec{g}_{0}\equiv(\boldsymbol{\mu}_{0},\mathbf{R}_{0},\mathbf{s},o,\mathbf{c})$, where $\boldsymbol{\mu}_{0}\in\mathbb{R}^{3}$, $\mathbf{R}_{0}\in\mathbb{SO}(3)$ are the 3D mean and orientation in the canonical frame, and $\mathbf{s}\in\mathbb{R}^{3}$ the scale, $o\in\mathbb{R}$ the opacity, and $\mathbf{c}\in\mathbb{R}^{3}$ the color, are persistent properties shared across time. To render a set of 3D Gaussians from a camera with world-to-camera extrinsics $\mathbf{E}$ and intrinsics $\mathbf{K}$, the projections of the 3D Gaussians in the image plane are obtained by 2D Gaussians parameterized as $\boldsymbol{\mu}^{\prime}_{0}\in\mathbb{R}^{2}$ and $\mathbf{\Sigma}^{\prime}_{0}\in\mathbb{R}^{2\times 2}$ via affine approximation
$\boldsymbol{\mu}^{\prime}_{0}(\mathbf{K},\mathbf{E})=\Pi(\mathbf{K}\mathbf{E}\boldsymbol{\mu}_{0})\in\mathbb{R}^{2},\quad\mathbb{\mathbf{\Sigma}}^{\prime}_{0}(\mathbf{K},\mathbf{E})=\mathbf{J}_{\mathbf{K}\mathbf{E}}\mathbf{\Sigma}_{0}\mathbf{J}_{\mathbf{K}\mathbf{E}}^{\top}\in\mathbb{R}^{2\times 2},$ (1)
@@ 3.2 Dynamic Scene Representation
Rasterizing 3D trajectories. Given this representation, we now describe how we obtain pixelwise 3D motion trajectory at any query frame $I_{t}$. we take a similar approach to Wang et al. [94] and rasterize the motion trajectories of 3D Gaussians into query frame $I_{t}$. Namely, for a query camera at time $t$ with intrinsics $\mathbf{K}_{t}$ and extrinsics $\mathbf{E}_{t}$, we perform rasterization to obtain a map ${}^{\textrm{w}}{\hat{\mathbf{X}}}_{t\rightarrow t^{\prime}}\in\mathcal{R}^{H\times W\times 3}$ that contains the expected 3D world coordinates of the surface points corresponding to each pixel at target time $t^{\prime}$
@@
$\hat{\mathbf{U}}_{t\rightarrow t^{\prime}}(\mathbf{p})=\Pi\big{(}\mathbf{K}_{t^{\prime}}{}^{\textrm{c}}\hat{\mathbf{X}}_{t\rightarrow t^{\prime}}(\mathbf{p})\big{)},~{}\hat{\mathbf{D}}_{t\rightarrow t^{\prime}}(\mathbf{p})=\left({}^{\textrm{c}}\hat{\mathbf{X}}_{t\rightarrow t^{\prime}}(\mathbf{p})\right)_{[3]}$ (6)

### v2

# Shape of Motion: 4D Reconstruction from a Single Video

@@ Abstract
Monocular dynamic reconstruction is a challenging and long-standing vision problem due to the highly ill-posed nature of the task. Existing approaches depend on templates, are effective only in quasi-static scenes, or fail to model 3D motion explicitly. We introduce a method for reconstructing generic dynamic scenes, featuring explicit, persistent 3D motion trajectories in the world coordinate frame, from casually captured monocular videos. We tackle the problem with two key insights: First, we exploit the low-dimensional structure of 3D motion by representing scene motion with a compact set of $\mathbb{SE}(3)$ motion bases. Each point’s motion is expressed as a linear combination of these bases, facilitating soft decomposition of the scene into multiple rigidly-moving groups. Second, we take advantage of off-the-shelf data-driven priors such as monocular depth maps and long-range 2D tracks, and devise a method to effectively consolidate these noisy supervisory signals, resulting in a globally consistent representation of the dynamic scene. Experiments show that our method achieves state-of-the-art performance for both long-range 3D/2D motion estimation and novel view synthesis on dynamic scenes. Project Page: https://shape-of-motion.github.io/ footnotetext: Equal contribution
Figure 1: Shape of Motion. Our method enables joint long-range 3D tracking and novel view synthesis from a monocular video of a complex dynamic scene. We render moving elements at a fixed viewpoint across time and visualize estimated 3D motion as colorful trajectories. These trajectories reveal distinct geometric patterns, which leads to the term “Shape of Motion”.
@@ 1 Introduction
Reconstructing the persistent geometry and their 3D motion across a video is crucial for understanding and interacting with the underlying physical world. While recent years have seen impressive progress in modeling static 3D scenes [71, 46], recovering the geometry and motion of complex dynamic 3D scenes, especially from a single video, remains an open challenge. A number of prior dynamic reconstruction and novel view synthesis approaches have attempted to tackle this problem. However, most methods rely on synchronized multi-view videos [9, 23, 23, 69, 111] or additional LIDAR/depth sensors [29, 101, 65, 80, 109]. Recent monocular approaches can operate on regular dynamic videos, but they typically model 3D scene motion as short-range scene flow between consecutive times [59, 60, 24] or deformation fields that map between canonical and view space [76, 77, 122], failing to capture 3D motion trajectories persistent over a video.
The longstanding challenge for general in-the-wild videos lies in the poorly constrained nature of the 4D reconstruction problem. In this work, we tackle this challenge with two key insights. The first is that, while the image space dynamics can be complex and discontinuous, the underlying 3D motion is a composition of continuous simple rigid motions. Our second insight is that data-driven priors provide complementary, though noisy cues, that aggregate well into a globally coherent representation of the 3D scene geometry and motion.
Motivated by these two insights, we represent the dynamic scene as a set of persistent 3D Gaussians, and represent their motion across the video in terms of a compact set of shared $\mathbb{SE}(3)$ motion bases. Unlike traditional scene flow, which computes 3D correspondences between consecutive frames, our representation recovers a persistent 3D trajectory over the whole video, enabling long-range 3D tracking across the entire video. As the 3D trajectories produced by our method capture the geometric patterns that trace each point’s movement through 3D space and time as shown in Figure 1, we refer to our approach as “Shape of Motion”. We show how to fit our explicit scene representation to a general video in-the-wild, by fusing together complementary cues from two main sources: monocular depth estimates per-frame, and 2D track estimates across frames. We conduct extensive evaluations on both synthetic and real-world dynamic video datasets, and show that our proposed approach significantly outperforms prior methods in both long-range 2D/3D tracking and novel view synthesis.
In summary, our key contributions are: (1) A 4D scene representation enabling both real-time novel view synthesis and globally consistent 3D tracking for any point at any time. (2) An optimization framework that fuses learned motion and geometry priors into a unified 4D representation from a single, casually captured monocular video.
@@ 2 Related Work
Correspondences and Tracking. While monocular 3D long range tracking remains largely unexplored, numerous approaches track in 2D image space, typically using optical flow for point correspondences. This involves estimating dense motion fields between image pairs [33, 4, 68, 8, 7, 96, 18, 36, 43, 97, 40, 36, 115, 44, 42, 82, 90]. Sparse keypoint matching methods can enable long trajectory generation [2, 84, 66, 14, 63], but these methods are primarily intended for sparse 3D reconstruction. Long-range 2D trajectory estimation for arbitrary points has been explored in earlier works, which relied on hand-crafted priors to generate motion trajectories [83, 3, 86, 87, 91, 102]. Recently, there has been a resurgence of interest in this problem, with several works showcasing impressive long-range 2D tracking results on challenging, in-the-wild videos. These approaches employ either test-time optimization where models consolidate noisy short-range motion estimates into long-term correspondences [104, 72, 93], or data-driven strategies [30, 45, 16, 17], where neural networks learn long-term correspondence estimates from synthetic data [15, 129]. While these methods effectively track any 2D point throughout a video, they lack the knowledge of underlying 3D scene geometry and motions.
To model 3D scene motion, recent work in monocular settings has explored self-supervised learning and test-time optimization [118, 38, 59, 60, 117, 119]. More recent approaches estimate 3D trajectories in general scenes in a feedforward manner [113, 74, 52], but the predicted motion remains in the frame space, entangling object and camera motion. In contrast, our method recovers persistent, long-range 3D trajectories in the world coordinate space.
Dynamic Reconstruction and View Synthesis. Our work also relates to dynamic 3D scene reconstruction and novel view synthesis. In non-rigid reconstruction, early methods often required RGB-D sensors [130, 73, 5, 19, 41] or strong hand-crafted priors [85, 54, 81]. A family of works [75, 13, 6] apply low-rank assumptions to the underlying motion model. Recent work has demonstrated progress toward general dynamic reconstruction either by integrating monocular depth priors [58, 127, 70, 51, 128, 62], or directly learning to perform dynamic reconstruction [106, 126, 67, 105] from data. However, these methods do not perform long-range 3D tracking, and do not focus on novel view synthesis.
@@
Neural Radiance Fields (NeRF) [71] and Gaussian Splatting [47] have shown strong performance in novel view synthesis. For dynamic scenes, many methods [1, 23, 9, 57, 23, 10, 92, 95, 103, 61] still require simultaneous multi-view video observations or predefined templates [39, 56, 110]. Template-free monocular approaches model dynamic scenes with different types of representations such as video depth maps [125], time-dependent NeRFs [20, 59, 79, 76, 100, 77, 112, 60], and dynamic 3D Gaussians [111, 21, 123, 122]. Although significant progress has been made, as DyCheck [25] pointed out, many approaches focus on scenarios with camera teleportation [53, 122] or quasi-static scenes [76, 77], which are effectively multi-view and do not represent real-world monocular videos. In this work, we focus on casually-captured monocular videos, a more practical and challenging setup. Recent and concurrent efforts [55, 107, 94, 64, 12] have also investigated similar settings, using strong off-the-shelf data-driven priors.
@@
Figure 2: System Overview. Given a single RGB video sequence with known camera poses, monocular depth maps and 2D tracks computed from off-the-shelf models [121, 16] as input, we optimize a dynamic scene representation as a set of persistent 3D Gaussians that translate and rotate over time. To capture the low-dimensional nature of scene motion, we model the motion with a set of compact $\mathbb{SE}(3)$ motion bases shared across all scene elements. Each 3D Gaussian’s motion is represented as a linear combination of these motion bases, weighted by motion coefficients specific to each Gaussian. We supervise our scene representation (canonical 3D Gaussians, per-Gaussian motion coefficients, and global motion bases) by comparing the rendered outputs (RGB, depths and 2D tracks) with the corresponding input signals. This results in a globally coherent dynamic 3D scene representation in the world coordinate with long-range 3D trajectories.
@@ 3 Method
Our method takes as input a sequence of $T$ video frames $\{I_{t}\in\mathbb{R}^{H\times W\times 3}\}$ of a dynamic scene, the camera intrinsics $\mathbf{K}_{t}\in\mathbb{R}^{3\times 3}$, and world-to-camera extrinsics $\mathbf{E}_{t}\in\mathbb{SE}(3)$ of each input frame $I_{t}$. From these inputs, we aim to recover the geometry of the entire dynamic scene and the full-length 3D motion trajectory of every point in the scene. Unlike most prior dynamic NeRFs methods [58, 60, 25, 112, 100] which render the scene contents through volumetric ray casting and represent the motion implicitly at fixed 3D locations, we model the dense scene elements as a set of canonical 3D Gaussians, that translate and rotate over entire video as persistent motion trajectories. We adopt explicit point-based representation because it simultaneously allows for both (1) high-fidelity rendering in real-time and (2) full-length 3D tracking of any surface point from any input time.
Optimizing an explicit representation of dynamic 3D Gaussians from a single video is severely ill-posed — at each point in time, the moving subjects in the scene are observed from only a single viewpoint. In order to overcome this ambiguity, we make two insights: First, while the projected 2D dynamics might be complex in the video, the underlying 3D motion in the scene is low-dimensional, and composed of simpler units of rigid motion. Second, powerful data-driven priors, namely monocular depth estimates and long-range 2D tracks, provide complementary but noisy signals of the underlying 3D scene. We propose a system that fuses these noisy estimates together into a globally coherent representation of both the scene geometry and motion. We show a schematic of our pipeline in Figure 2.
@@ 3.1 Preliminaries: 3D Gaussian Splatting
We represent the appearance and geometry of a dynamic scene with a global set of 3D Gaussians, an explicit and expressive differentiable scene representation [46] for efficient optimization and rendering. We define parameters of each 3D Gaussian in the canonical frame $t_{0}$ as $\vec{g}_{0}\equiv(\boldsymbol{\mu}_{0},\mathbf{R}_{0},\mathbf{s},o,\mathbf{c})$, where $\boldsymbol{\mu}_{0}\in\mathbb{R}^{3}$, $\mathbf{R}_{0}\in\mathbb{SO}(3)$ are the 3D mean and orientation in the canonical frame, and $\mathbf{s}\in\mathbb{R}^{3}$ the scale, $o\in\mathbb{R}$ the opacity, and $\mathbf{c}\in\mathbb{R}^{3}$ the color. $\Sigma_{0}$ is the canonical frame covariance matrix. These properties are persistent and shared across time. To render a set of 3D Gaussians from a camera with world-to-camera extrinsics $\mathbf{E}$ and intrinsics $\mathbf{K}$, the projections of the 3D Gaussians in the image plane are obtained by 2D Gaussians parameterized as $\boldsymbol{\mu}^{\prime}_{0}\in\mathbb{R}^{2}$ and $\mathbf{\Sigma}^{\prime}_{0}\in\mathbb{R}^{2\times 2}$ via affine approximation
$\boldsymbol{\mu}^{\prime}_{0}(\mathbf{K},\mathbf{E})$ $=\Pi(\mathbf{K}\mathbf{E}\boldsymbol{\mu}_{0})\in\mathbb{R}^{2},$
$\mathbb{\mathbf{\Sigma}}^{\prime}_{0}(\mathbf{K},\mathbf{E})$ $=\mathbf{J}_{\mathbf{K}\mathbf{E}}\mathbf{\Sigma}_{0}\mathbf{J}_{\mathbf{K}\mathbf{E}}^{\top}\in\mathbb{R}^{2\times 2}\textyen,$ (1)
@@ 3.2 Dynamic Scene Representation
Rasterizing 3D Trajectories. Given this representation, we now describe how we obtain pixelwise 3D motion trajectory at any query frame $I_{t}$. we take a similar approach to Wang et al[104] and rasterize the motion trajectories of 3D Gaussians into query frame $I_{t}$. Namely, for a query camera at time $t$ with intrinsics $\mathbf{K}_{t}$ and extrinsics $\mathbf{E}_{t}$, we perform rasterization to obtain a map $^{\textrm{w}}{\hat{\mathbf{X}}}_{t\rightarrow t^{\prime}}\in\mathcal{R}^{H\times W\times 3}$ that contains the expected 3D world coordinates of the surface points corresponding to each pixel at target time $t^{\prime}$
@@
$\hat{\mathbf{U}}_{t\rightarrow t^{\prime}}(\mathbf{p})=\Pi\big(\mathbf{K}_{t^{\prime}}{}^{\textrm{c}}\hat{\mathbf{X}}_{t\rightarrow t^{\prime}}(\mathbf{p})\big),~\hat{\mathbf{D}}_{t\rightarrow t^{\prime}}(\mathbf{p})=\left({}^{\textrm{c}}\hat{\mathbf{X}}_{t\rightarrow t^{\prime}}(\mathbf{p})\right)_{[3]},$ (6)